OpenStudy (anonymous):

The school that Willie goes to is selling tickets to a school performance. On the first day of ticket sales the school sold 3 senior citizen tickets and 12 student tickets for a total of $213. The school took in $85 on the second day by selling 5 senior citizen tickets and 2 student tickets. What is the price each of one senior citizen and one student ticket?

OpenStudy (radar):

Let x be the price of senior tickets Let y be the price of student tickets Day 1. 3x+12y=$213.00 Day 2. 5x+2y=$85.00 Take day 2 equation and multiply all by 6 so that Day 2. 30x+12y=$510.00 now subtract Day 1 equation Day 1 -3x-12y=-213.00 getting 27x = 297 divide by 27 getting x=$11.00 Senior tickets cost $11.00 substitute in day 1 and find y the students ticket cost 3(11.00)+12y=213.00 33.00 +12y =213.00 subtract 33 from both sides 12y=213.00-33=180.00 divide by 12 y=$15.00 Students ticket cost $15.00 each.

OpenStudy (radar):

Day 2 equation was multiplied by 6 for the sole reason of getting the y variable to equal 12, so when subtracted from day 1 equation the y value would drop out leaving only x so it could be solved.